# 0-supple.tex

### v1

@@ Appendix B Statistical Inference in the Distributed Environment
3:     Compute a local estimator
4:end for
6:for $j=1,2,\dots,J$ do
8:end for
Appendix C Technical Proof of the Theoretical Results
C.1 Proof of the Results for the Homogeneous Setting
@@
For $j\in[J]$, $\ell\in[L]$, define $Z_{j,\ell}=\mathcal{M}_{j}(\mathcal{Z}_{\ell})$, $T_{j}=\mathcal{M}_{j}(\mathcal{T})$, and $G_{j}=\mathcal{M}_{j}(\mathcal{G})$. Let $U_{j\perp}\in\mathbb{R}^{p_{j}\times(p_{j}-r_{j})}$ be the orthogonal complement of $U_{j}$, i.e., $[U_{j}\,\,U_{j\perp}]$ is an orthogonal matrix in $\mathbb{R}^{p_{j}\times p_{j}}$. Denote the compact singular value decomposition of $G_{j}$ by $U_{G_{j}}\Lambda_{j}V^{\top}_{G_{j}}$, where $U_{G_{j}}\in\mathbb{O}^{r_{j}\times r_{j}}$, $\Lambda_{j}\in\mathbb{R}^{r_{j}\times r_{j}}$, and $V^{\top}_{G_{j}}V_{G_{j}}=I_{r_{j}}$. Note the model (1) is equivalent to the model with $\widetilde{\mathcal{G}}=\mathcal{G}\times_{1}O^{\top}_{1}\times_{2}O^{\top}_{2}\cdots\times_{J}O^{\top}_{J}$ and $\widetilde{U}_{j}=U_{j}O_{j}$, for any $O_{j}\in\mathbb{O}^{r_{j}\times r_{j}}$, $j\in[J]$. Taking $O_{j}=U_{G_{j}}$ leads to $\widetilde{G}_{j}=\Lambda_{j}V_{G_{j}}^{\top}$ and thus $\widetilde{G}_{j}\widetilde{G}_{j}^{\top}=\Lambda_{j}^{2}$, while the projection matrix $U_{j}O_{j}O_{j}^{\top}U_{j}^{\top}=U_{j}U_{j}^{\top}$ remains invariant. Therefore, without loss of generality, we assume $G_{j}G_{j}^{\top}=\Lambda_{j}^{2}$ in the sequel. Moreover, by the definition of $\lambda_{\min}$ and $\kappa_{0}$, it holds that
@@
For each $j\in[J]$, $\ell\in[L]$, define a “locally-good” event:
@@
and a “globally-good” event
$E(C):=\bigcap_{\ell=1}^{L}\bigcap_{j=1}^{J}E_{j,\ell}(C).$
We have the following lemma for the probability of $E(C)$.
Lemma C.1.
Under the assumptions in Theorem 2.1, there exist constant $C_{1}$, $c_{1}$, $C_{2}$ such that $\mathbb{P}[E(C_{2})]\geq 1-C_{1}Le^{-c_{1}p}$.
Note that we assume that $L\lesssim p^{c_{3}}$ for some $c_{3}>0$, which implies that $E(C_{2})$ has probability approaching one. For the homogeneous setting, we follow Xia et al. (2022) to decompose $\widehat{U}_{j}\widehat{U}_{j}^{\top}-U_{j}U_{j}^{\top}$. By definition, the local estimator $\widehat{U}_{j,\ell}$ is composed of the first $r_{j}$ left singular vectors of
$\mathcal{M}_{j}\big{(}\mathcal{T}_{\ell}\times_{1}\widehat{U}_{1,\ell}^{(0)\top}\times_{2}\widehat{U}_{2,\ell}^{(0)\top}\cdots\times_{j-1}\widehat{U}_{j-1,\ell}^{(0)\top}\times_{j+1}\widehat{U}_{j+1,\ell}^{(0)\top}\cdots\times_{J}\widehat{U}_{J,\ell}^{(0)\top}\big{)}=(T_{j}+Z_{j,\ell})\widehat{U}^{(0)}_{-j},$
@@
$\quad(T_{j}+Z_{j,\ell})\widehat{U}^{(0)}_{-j}\widehat{U}_{-j}^{(0)\top}(T^{\top}_{j}+Z^{\top}_{j,\ell})=T_{j}U_{-j}U_{-j}^{\top}T_{j}^{\top}+\mathfrak{E}_{j,\ell}=U_{j}\Lambda_{j}^{2}U_{j}^{\top}+\mathfrak{E}_{j,\ell},$ (20)
@@
$\widehat{U}^{(0)}_{-j}=\widehat{U}^{(0)}_{1}\otimes\widehat{U}^{(0)}_{2}\otimes\cdots\otimes\widehat{U}^{(0)}_{j-1}\otimes\widehat{U}^{(0)}_{j+1}\otimes\cdots\otimes\widehat{U}^{(0)}_{J},$
@@
$\zeta_{j,\ell,2}$ $:=T_{j}\left[\widehat{U}^{(0)}_{-j}\widehat{U}_{-j}^{(0)\top}-U_{-j}U_{-j}^{\top}\right]Z^{\top}_{j,\ell},$
@@
$\zeta_{j,\ell,4}$ $:=Z_{j,\ell}\left[\widehat{U}^{(0)}_{-j}\widehat{U}_{-j}^{(0)\top}-U_{-j}U_{-j}^{\top}\right]Z_{j,\ell}^{\top},$
$\zeta_{j,\ell,5}$ $:=T_{j}\left[\widehat{U}^{(0)}_{-j}\widehat{U}_{-j}^{(0)\top}-U_{-j}U_{-j}^{\top}\right]T_{j}^{\top}.$
@@
Lemma C.2.
Under the event $E(C_{2})$ and the assumptions in Theorem 2.1, there exists some absolute constant $C_{2}^{\prime}>0$ such that
$\left\lVert\zeta_{j,\ell,1}\right\rVert_{2}\leq C_{2}^{\prime}\kappa_{0}\lambda_{\min}\sigma\sqrt{p},\quad\left\lVert\zeta_{j,\ell,2}\right\rVert_{2}\leq C_{2}^{\prime}\kappa_{0}p\sigma^{2}\sqrt{r},$ (22)
$\left\lVert\zeta_{j,\ell,3}\right\rVert_{2}\leq C_{2}^{\prime}p\sigma^{2},\quad\left\lVert\zeta_{j,\ell,4}\right\rVert_{2}\leq C_{2}^{\prime}p^{3/2}\sqrt{r}\sigma^{3}\lambda_{\min}^{-1},\quad\left\lVert\zeta_{j,\ell,5}\right\rVert_{2}\leq C_{2}^{\prime}\kappa_{0}^{2}p\sigma^{2},$
By Lemma C.2, under the event $E(C_{2})$, $\left\lVert\mathfrak{E}_{j,\ell}\right\rVert_{2}\lesssim\kappa_{0}\lambda_{\min}\sigma\sqrt{p}<\lambda^{2}_{\min}/2$. Applying Theorem 1 in Xia (2021) yields
@@
$\quad-U_{j\perp}U_{j\perp}^{\top}\mathfrak{E}^{(1)}_{j,\ell}U_{j}\Lambda_{j}^{-2}U_{j}^{\top}\mathfrak{E}_{j,\ell}U_{j}\Lambda_{j}^{-2}U_{j}^{\top}+\mathfrak{R}_{j,\ell},$
@@
$\mathfrak{S}_{j,\ell,2}$ $:=U_{j}\Lambda_{j}^{-2}U_{j}^{\top}\zeta^{(t)}_{j,\ell,2}U_{j\perp}U_{j\perp}^{\top}+U_{j\perp}U_{j\perp}^{\top}\zeta^{\top}_{j,\ell,2}U_{j}\Lambda_{j}^{-2}U_{j}^{\top}$
@@
$\quad-U_{j\perp}U_{j\perp}^{\top}\zeta^{\top}_{j,\ell,1}U_{j}\Lambda_{j}^{-2}U_{j}^{\top}(\zeta_{j,\ell,1}+\zeta^{\top}_{j,\ell,1})U_{j}\Lambda_{j}^{-2}U_{j}^{\top},$
and
$\mathfrak{S}_{j,\ell,3}$ $:=U_{j}\Lambda_{j}^{-2}U_{j}^{\top}\zeta_{j,\ell,4}U_{j\perp}U_{j\perp}^{\top}+U_{j\perp}U_{j\perp}^{\top}\zeta_{j,\ell,4}U_{j}\Lambda_{j}^{-2}U_{j}^{\top}$ (26)
$\quad-U_{j}\Lambda_{j}^{-2}U_{j}^{\top}\left[\mathfrak{E}_{j,\ell}-\zeta_{j,\ell,1}-\zeta_{j,\ell,1}^{\top}\right]U_{j}\Lambda_{j}^{-2}U_{j}^{\top}\mathfrak{E}_{j,\ell}U_{j\perp}U_{j\perp}^{\top}$
$\quad-U_{j}\Lambda_{j}^{-2}U_{j}^{\top}\mathfrak{E}_{j,\ell}U_{j}\Lambda_{j}^{-2}U_{j}^{\top}\left[\mathfrak{E}_{j,\ell}-\zeta_{j,\ell,1}-\zeta_{j,\ell,1}^{\top}\right]U_{j\perp}U_{j\perp}^{\top}$
$\quad-U_{j}\Lambda_{j}^{-2}U_{j}^{\top}\left[\mathfrak{E}_{j,\ell}-\zeta_{j,\ell,1}-\zeta_{j,\ell,1}^{\top}\right]U_{j\perp}U_{j\perp}^{\top}\mathfrak{E}_{j,\ell}U_{j}\Lambda_{j}^{-2}U_{j}^{\top}$
$\quad-U_{j}\Lambda_{j}^{-2}U_{j}^{\top}\mathfrak{E}_{j,\ell}U_{j\perp}U_{j\perp}^{\top}\left[\mathfrak{E}_{j,\ell}-\zeta_{j,\ell,1}-\zeta_{j,\ell,1}^{\top}\right]U_{j}\Lambda_{j}^{-2}U_{j}^{\top}$
$\quad-U_{j\perp}U_{j\perp}^{\top}\left[\mathfrak{E}_{j,\ell}-\zeta_{j,\ell,1}-\zeta_{j,\ell,1}^{\top}\right]U_{j}\Lambda_{j}^{-2}U_{j}^{\top}\mathfrak{E}_{j,\ell}U_{j}\Lambda_{j}^{-2}U_{j}^{\top}$
$\quad-U_{j\perp}U_{j\perp}^{\top}\mathfrak{E}_{j,\ell}U_{j}\Lambda_{j}^{-2}U_{j}^{\top}\left[\mathfrak{E}_{j,\ell}-\zeta_{j,\ell,1}-\zeta_{j,\ell,1}^{\top}\right]U_{j}\Lambda_{j}^{-2}U_{j}^{\top}+\mathfrak{R}_{j,\ell}.$
Under $E(C_{2})$, it holds that $\left\lVert Z_{j,\ell}U_{-j}\right\rVert_{2}\leq C_{2}\sigma\sqrt{p}$. Using the fact that $\left\lVert\Lambda_{j}^{-1}G_{j}\right\rVert_{2}=1$ and $\left\lVert\Lambda_{j}^{-1}\right\rVert_{2}\leq\lambda_{\min}^{-1}$, we have
$\left\lVert\mathfrak{S}_{j,\ell,1}\right\rVert_{2}=\left\lVert U_{j}\Lambda_{j}^{-2}G_{j}U_{-j}^{\top}Z_{j,\ell}^{\top}U_{j\perp}U_{j\perp}^{\top}+U_{j\perp}U_{j\perp}^{\top}Z_{j,\ell}U_{-j}G_{j}^{\top}\Lambda_{j}^{-2}U_{j}^{\top}\right\rVert_{2}\lesssim\lambda^{-1}_{\min}\sigma\sqrt{p},$ (27)
with probability at least $1-C_{1}e^{-c_{1}p}$. By (22), it holds that
@@
Since $Z_{j,\ell}$ has i.i.d. $\mathcal{N}(0,\sigma^{2})$ entries, the averaged noise $\overline{Z}_{j}$ has i.i.d. $\mathcal{N}\left(0,\frac{\sigma^{2}}{L}\right)$ entries. By the proof of Lemma C.1, it holds that $\left\lVert\overline{Z}_{j}U_{-j}\right\rVert_{2}\lesssim\sigma\sqrt{p/L}$ and thus
$\left\lVert\overline{\mathfrak{S}_{j,1}}\right\rVert_{2}=\left\lVert U_{j}\Lambda_{j}^{-2}G_{j}U_{-j}^{\top}\overline{Z}^{\top}_{j}U_{j\perp}U_{j\perp}^{\top}+U_{j\perp}U_{j\perp}^{\top}\overline{Z}_{j}U_{-j}G_{j}^{\top}\Lambda_{j}^{-2}U_{j}^{\top}\right\rVert_{2}\lesssim\lambda^{-1}_{\min}\sigma\sqrt{p/L},$ (30)
with probability at least $1-C_{1}e^{-c_{1}p}$.

### v3

@@ Appendix C Further Results for Knowledge Transfer
3:     for $j=1,2,\dots,J$ do
5:         Send $\widehat{U}_{j,\ell}$ to the target machine;
6:     end for
7:end for
8:for $j=1,2,\dots,J$ do
9:     On the target machine, compute $\widehat{U}_{j}=\mathrm{svd}_{r_{j,U}}\Big{[}\sum_{\ell=1}^{L}w_{\ell}\widehat{U}_{j,\ell}\widehat{U}^{\top}_{j,\ell}\Big{]}$;
10:     Compute $\widehat{V}_{j,1}=\mathrm{svd}_{r_{j,V,1}}\Big{[}\big{(}I_{p_{j}}-\widehat{U}_{j}\widehat{U}_{j}^{\top}\big{)}\widetilde{M}_{j,1}\Big{]}$;
11:end for
To transfer knowledge from $\mathcal{T}_{2},\dots,\mathcal{T}_{L}$ to improve the learning performance on $\mathcal{T}_{1}$, we first generalize Algorithm 3 to Algorithm 4 and then provide the theoretical guarantee for the estimators. Similar to Section 4, for $j\in[J]$ and $\ell\in[L]$, let $\Lambda_{j,\ell}$ be the $r_{j,\ell}\times r_{j,\ell}$ singular value matrix of $\mathcal{M}_{j}(\mathcal{G}_{\ell})$. Define $\lambda_{\max}$, $\lambda_{\min}$ to be the maximum and minimum singular value over all $\Lambda_{j,\ell}$, respectively, and let $\kappa_{0}=\lambda_{\max}\lambda^{-1}_{\min}$. Moreover, define $\Delta=\min_{j\in[J],\ell\in[L]}\{\lambda_{r_{j,U},j,\ell}-\lambda_{r_{j,U}+1,j,\ell}\}$ and $\kappa=\lambda_{\max}/\Delta$, where $\lambda_{r,j,\ell}$ denotes the $r$-th largest singular value of $\Lambda_{j,\ell}$. Additionally, let $r=\max_{j,\ell}r_{j,\ell}$ and $r_{V}=\max_{j,\ell}r_{j,V,\ell}$.
Theorem C.1.
Assume that there exist constants $C_{1},c_{1},C_{2}$ such that, with probability at least $1-C_{1}e^{-c_{1}p}$, $\big{\|}\widehat{U}_{j,\ell}^{(0)}\widehat{U}_{j,\ell}^{(0)\top}+\widehat{V}_{j,\ell}^{(0)}\widehat{V}_{j,\ell}^{(0)\top}-U_{j}U_{j}^{\top}-V_{j,\ell}V_{j,\ell}^{\top}\big{\|}_{2}\leq C_{2}\sqrt{p}\sigma_{\ell}\lambda_{\min}^{-1}$ for all $j\in[J],\ell\in[L]$. Assume $p_{j}\asymp p$ for all $j$, $r^{J-1}=O(p)$, $\kappa_{0}=O(1)$, $\kappa=O(1)$, and $\sqrt{prL}=o(\min(\Delta,\lambda_{\min})/\max_{\ell}\sigma_{\ell})$. Then there exist constants $\widetilde{C}_{1},\widetilde{c}_{1},\widetilde{C}_{2}$ such that
$\sup_{j}\left\lVert\widehat{U}_{j}\widehat{U}_{j}^{\top}-U_{j}U_{j}^{\top}\right\rVert_{\rm F}$ $\leq\widetilde{C}_{2}\frac{\sqrt{pr}\sqrt{\sum_{\ell=1}^{L}w_{\ell}^{2}\sigma_{\ell}^{2}}}{\Delta},$ (17)
$\sup_{j}\left\lVert\widehat{V}_{j,1}\widehat{V}_{j,1}^{\top}-V_{j,1}V_{j,1}^{\top}\right\rVert_{\rm F}$ $\leq\widetilde{C}_{2}\left(\frac{\sqrt{pr_{V}}}{\lambda_{\min}}\Big{(}\sigma_{1}+\sqrt{r}\sqrt{\sum_{\ell=1}^{L}w_{\ell}^{2}\sigma_{\ell}^{2}}\Big{)}\right),$ (18)
with probability at least $1-\widetilde{C}_{1}e^{-\widetilde{c}_{1}p}$, where $\widehat{U}_{j},\widehat{V}_{j,1}$’s are the outputs of Algorithm 4.
Theorem C.1 is a generalization of Theorem 4.1 from $L=2$ to arbitrary $L$. Furthermore, it also broadens the results for the heterogeneous setting established in Section 3 to accommodate varying noise levels across different tensors $\mathcal{T}_{\ell}$. If $\sigma_{\ell}=\sigma$ for all $\ell$ and $w_{\ell}=\frac{1}{L}$, the results of Theorem C.1 reduce to those of Theorem 3.1. The proof of Theorem C.1 is provided in Section D.3.
Based on the error rates established in (17) and (18), we further give the optimal choice for $w_{\ell}$ by minimizing $\sum_{\ell=1}^{L}w_{\ell}^{2}\sigma^{2}_{\ell}$ under the constraint $\sum_{\ell=1}^{L}w_{\ell}=1$, which yields the optimal weights $w_{\ell}^{*}=\frac{\sigma_{\ell}^{-2}}{\sum_{\ell=1}^{L}\sigma_{\ell}^{-2}}$. The corresponding optimal error rates are
$\sup_{j}\left\lVert\widehat{U}_{j}\widehat{U}_{j}^{\top}-U_{j}U_{j}^{\top}\right\rVert_{\rm F}=O_{\mathbb{P}}\left(\frac{\sqrt{pr}\overline{\sigma}}{\Delta}\right)\,\mathrm{and}\,\sup_{j}\left\lVert\widehat{V}_{j,1}\widehat{V}_{j,1}^{\top}-V_{j,1}V_{j,1}^{\top}\right\rVert_{\rm F}=O_{\mathbb{P}}\left(\frac{\sqrt{pr_{V}}}{\lambda_{\min}}\big{(}\sigma_{1}+\sqrt{r}\overline{\sigma}\big{)}\right),$ (19)
where $\overline{\sigma}^{2}=1/\sum_{\ell=1}^{L}\sigma_{\ell}^{-2}$.
In practice, we can estimate $\sigma_{\ell}$ by
$\widehat{\sigma}_{\ell}=\Big{\|}\mathcal{T}_{\ell}-\mathcal{T}_{\ell}\times_{1}\big{[}\widehat{U}_{1,\ell}\;\widehat{V}_{1,\ell}\big{]}\big{[}\widehat{U}_{1,\ell}\;\widehat{V}_{1,\ell}\big{]}^{\top}\times_{2}\cdots\times_{J}\big{[}\widehat{U}_{J,\ell}\;\widehat{V}_{J,\ell}\big{]}\big{[}\widehat{U}_{J,\ell}\;\widehat{V}_{J,\ell}\big{]}^{\top}\Big{\|}_{\rm F}/\sqrt{p_{1}p_{2}\cdots p_{J}},\quad\ell\in[L],$ (20)
and choose $\widehat{w}_{\ell}=\frac{\widehat{\sigma}_{\ell}^{-2}}{\sum_{\ell=1}^{L}\widehat{\sigma}_{\ell}^{-2}}$. We further provide the theoretical guarantee for the estimators obtained in Algorithm 4 with data-adaptive weights $\widehat{w}_{\ell}$.
Theorem C.2.
Let $\widehat{U}_{j},\widehat{V}_{j,1}$’s be the outputs of Algorithm 4 with weights $\widehat{w}_{\ell}$ defined above. Under the conditions of Theorem C.1, we have that
$\sup_{j}\left\lVert\widehat{U}_{j}\widehat{U}_{j}^{\top}-U_{j}U_{j}^{\top}\right\rVert_{\rm F}\leq\widetilde{C}_{2}\bigg{(}\frac{\sqrt{pr}\overline{\sigma}}{\Delta}+\frac{r\sqrt{L}\overline{\sigma}}{p^{J/2-1}\Delta}\bigg{)},$ (21)
and
$\sup_{j}\left\lVert\widehat{V}_{j,1}\widehat{V}_{j,1}^{\top}-V_{j,1}V_{j,1}^{\top}\right\rVert_{\rm F}\leq\widetilde{C}_{2}\frac{\sqrt{pr_{V}}}{\lambda_{\min}}\bigg{(}\sigma_{1}+\sqrt{r}\overline{\sigma}+\frac{r\sqrt{L}\overline{\sigma}}{p^{(J-1)/2}}\bigg{)},$ (22)
for some constants $\widetilde{C}_{1}$ and $\widetilde{C}_{2}$, with probability at least $1-\widetilde{C}_{1}p^{-1}$. If we further assume that $rL/p^{J-1}=o(1)$, then, with probability at least $1-\widetilde{C}_{1}p^{-1}$,
$\sup_{j}\left\lVert\widehat{U}_{j}\widehat{U}_{j}^{\top}-U_{j}U_{j}^{\top}\right\rVert_{\rm F}\leq\widetilde{C}_{2}\frac{\sqrt{pr}\overline{\sigma}}{\Delta},\quad\sup_{j}\left\lVert\widehat{V}_{j,1}\widehat{V}_{j,1}^{\top}-V_{j,1}V_{j,1}^{\top}\right\rVert_{\rm F}\leq\widetilde{C}_{2}\frac{\sqrt{pr_{V}}}{\lambda_{\min}}\big{(}\sigma_{1}+\sqrt{r}\overline{\sigma}\big{)},$
which matches the error rates established in (19).
The proof of Theorem C.2 is provided in Section D.3. Compared to (19), the error rate of $\widehat{U}_{j}$ in (21) contains an additional term $r\sqrt{L}\overline{\sigma}/(p^{J/2-1}\Delta)$ due to the estimation error between $\widehat{w}_{\ell}$ and $w_{\ell}^{*}$, which is dominated by the first term $\sqrt{pr}\overline{\sigma}/\Delta$ under a mild condition that $rL/p^{J-1}=o(1)$. The same is true for the error rate of $\widehat{V}_{j,1}$ in (22).
Appendix D Technical Proof of the Theoretical Results
D.1 Proof of the Results for the Homogeneous Setting
@@
For $j\in[J]$, $\ell\in[L]$, define $Z_{j,\ell}=\mathcal{M}_{j}(\mathcal{Z}_{\ell})$, $T_{j}=\mathcal{M}_{j}(\mathcal{T}^{*})$, and $G_{j}=\mathcal{M}_{j}(\mathcal{G})$. Let $U_{j\perp}\in\mathbb{R}^{p_{j}\times(p_{j}-r_{j})}$ be the orthogonal complement of $U_{j}$, i.e., $[U_{j}\,\,U_{j\perp}]$ is an orthogonal matrix in $\mathbb{R}^{p_{j}\times p_{j}}$. Denote the compact singular value decomposition of $G_{j}$ by $U_{G_{j}}\Lambda_{j}V^{\top}_{G_{j}}$, where $U_{G_{j}}\in\mathbb{O}^{r_{j}\times r_{j}}$, $\Lambda_{j}\in\mathbb{R}^{r_{j}\times r_{j}}$, and $V^{\top}_{G_{j}}V_{G_{j}}=I_{r_{j}}$. Note the model (1) is equivalent to the model with $\widetilde{\mathcal{G}}=\mathcal{G}\times_{1}O^{\top}_{1}\times_{2}O^{\top}_{2}\cdots\times_{J}O^{\top}_{J}$ and $\widetilde{U}_{j}=U_{j}O_{j}$, for any $O_{j}\in\mathbb{O}^{r_{j}\times r_{j}}$, $j\in[J]$. Taking $O_{j}=U_{G_{j}}$ leads to $\widetilde{G}_{j}=\Lambda_{j}V_{G_{j}}^{\top}$ and thus $\widetilde{G}_{j}\widetilde{G}_{j}^{\top}=\Lambda_{j}^{2}$, while the projection matrix $U_{j}O_{j}O_{j}^{\top}U_{j}^{\top}=U_{j}U_{j}^{\top}$ remains invariant. Therefore, without loss of generality, we assume $G_{j}G_{j}^{\top}=\Lambda_{j}^{2}$ in the sequel. Moreover, by the definition of $\lambda_{\min}$ and $\kappa_{0}$, it holds that
@@
For each $j\in[J]$, $\ell\in[L]$, define a “locally-good” event
@@
$\quad\sup_{\begin{subarray}{c}X_{j^{\prime}}\in\mathbb{R}^{p_{j^{\prime}}\times r_{j^{\prime}}}\\ \left\lVert X_{j^{\prime}}\right\rVert_{2}\leq 1,\forall j^{\prime}\in[J]\setminus\{j\}\end{subarray}}\left\lVert Z_{j,\ell}(X_{1}\otimes X_{2}\otimes\cdots\otimes X_{j-1}\otimes X_{j+1}\otimes\cdots\otimes X_{J})\right\rVert_{2}\leq C\sigma\sqrt{pr}\bigg{\}}.$
Then define a “globally-good” event
$E(C):=\bigcap_{j=1}^{J}\bigg{\{}\Big{\{}\bigcap_{\ell=1}^{L}E_{j,\ell}(C)\Big{\}}\bigcap\Big{\{}\left\lVert\overline{Z}_{j}U_{-j}\right\rVert_{2}\leq C\sigma\sqrt{p/L}\Big{\}}\bigg{\}},$
where $\overline{Z}_{j}:=\frac{1}{L}\sum_{\ell=1}^{L}Z_{j,\ell}$. We have the following lemma for the probability of $E(C)$.
Lemma D.1.
Under the assumptions in Theorem 2.1, there exist constants $C_{1}^{\prime}$, $c_{1}^{\prime}$, $C_{2}^{\prime}$ such that $\mathbb{P}[E(C_{2}^{\prime})]\geq 1-C_{1}^{\prime}Le^{-c_{1}^{\prime}p}$.
The proof of Lemma D.1 is provided in Section D.4. Note that we assume that $L\lesssim p^{c_{3}}$ for some $c_{3}>0$, which implies that $E(C_{2}^{\prime})$ has probability approaching one. For the homogeneous setting, we follow Xia et al. (2022) to decompose $\widehat{U}_{j}\widehat{U}_{j}^{\top}-U_{j}U_{j}^{\top}$. By definition, the local estimator $\widehat{U}_{j,\ell}$ is composed of the first $r_{j}$ left singular vectors of
$\mathcal{M}_{j}\big{(}\mathcal{T}_{\ell}\times_{1}\widehat{U}_{1,\ell}^{(0)\top}\times_{2}\widehat{U}_{2,\ell}^{(0)\top}\cdots\times_{j-1}\widehat{U}_{j-1,\ell}^{(0)\top}\times_{j+1}\widehat{U}_{j+1,\ell}^{(0)\top}\cdots\times_{J}\widehat{U}_{J,\ell}^{(0)\top}\big{)}=(T_{j}+Z_{j,\ell})\widehat{U}^{(0)}_{-j,\ell},$
@@
$\quad(T_{j}+Z_{j,\ell})\widehat{U}^{(0)}_{-j,\ell}\widehat{U}_{-j,\ell}^{(0)\top}(T^{\top}_{j}+Z^{\top}_{j,\ell})=T_{j}U_{-j}U_{-j}^{\top}T_{j}^{\top}+\mathfrak{E}_{j,\ell}=U_{j}\Lambda_{j}^{2}U_{j}^{\top}+\mathfrak{E}_{j,\ell},$ (25)
@@
$\widehat{U}^{(0)}_{-j,\ell}=\widehat{U}^{(0)}_{1,\ell}\otimes\widehat{U}^{(0)}_{2,\ell}\otimes\cdots\otimes\widehat{U}^{(0)}_{j-1,\ell}\otimes\widehat{U}^{(0)}_{j+1,\ell}\otimes\cdots\otimes\widehat{U}^{(0)}_{J,\ell},$
@@
$\zeta_{j,\ell,2}$ $:=T_{j}\left[\widehat{U}^{(0)}_{-j,\ell}\widehat{U}_{-j,\ell}^{(0)\top}-U_{-j}U_{-j}^{\top}\right]Z^{\top}_{j,\ell},$
@@
$\zeta_{j,\ell,4}$ $:=Z_{j,\ell}\left[\widehat{U}^{(0)}_{-j,\ell}\widehat{U}_{-j,\ell}^{(0)\top}-U_{-j}U_{-j}^{\top}\right]Z_{j,\ell}^{\top},$
$\zeta_{j,\ell,5}$ $:=T_{j}\left[\widehat{U}^{(0)}_{-j,\ell}\widehat{U}_{-j,\ell}^{(0)\top}-U_{-j}U_{-j}^{\top}\right]T_{j}^{\top}.$
@@
Lemma D.2.
Under the event $E(C_{2}^{\prime})$ and the assumptions in Theorem 2.1, there exists some absolute constant $C_{3}>0$ such that
$\left\lVert\zeta_{j,\ell,1}\right\rVert_{2}\leq C_{3}\kappa_{0}\lambda_{\min}\sigma\sqrt{p},\quad\left\lVert\zeta_{j,\ell,2}\right\rVert_{2}\leq C_{3}\kappa_{0}p\sigma^{2}\sqrt{r},$ (27)
$\left\lVert\zeta_{j,\ell,3}\right\rVert_{2}\leq C_{3}p\sigma^{2},\quad\left\lVert\zeta_{j,\ell,4}\right\rVert_{2}\leq C_{3}p^{3/2}\sqrt{r}\sigma^{3}\lambda_{\min}^{-1},\quad\left\lVert\zeta_{j,\ell,5}\right\rVert_{2}\leq C_{3}\kappa_{0}^{2}p\sigma^{2},$
By Lemma D.2, under the event $E(C_{2}^{\prime})$, $\left\lVert\mathfrak{E}_{j,\ell}\right\rVert_{2}\lesssim\kappa_{0}\lambda_{\min}\sigma\sqrt{p}<\lambda^{2}_{\min}/2$. Applying Theorem 1 in Xia (2021) yields
@@
$\quad-U_{j\perp}U_{j\perp}^{\top}\mathfrak{E}_{j,\ell}U_{j}\Lambda_{j}^{-2}U_{j}^{\top}\mathfrak{E}_{j,\ell}U_{j}\Lambda_{j}^{-2}U_{j}^{\top}+U_{j}\Lambda_{j}^{-4}U_{j}^{\top}\mathfrak{E}_{j,\ell}U_{j\perp}U_{j\perp}^{\top}\mathfrak{E}_{j,\ell}U_{j\perp}U_{j\perp}^{\top}$
$\quad+U_{j\perp}U_{j\perp}^{\top}\mathfrak{E}_{j,\ell}U_{j\perp}U_{j\perp}^{\top}\mathfrak{E}_{j,\ell}U_{j}\Lambda_{j}^{-4}U_{j}^{\top}+U_{j\perp}U_{j\perp}^{\top}\mathfrak{E}_{j,\ell}U_{j}\Lambda_{j}^{-4}U_{j}^{\top}\mathfrak{E}_{j,\ell}U_{j\perp}U_{j\perp}^{\top}+\mathfrak{R}_{j,\ell},$
@@
$\mathfrak{S}_{j,\ell,2}$ $:=U_{j}\Lambda_{j}^{-2}U_{j}^{\top}\zeta_{j,\ell,2}U_{j\perp}U_{j\perp}^{\top}+U_{j\perp}U_{j\perp}^{\top}\zeta^{\top}_{j,\ell,2}U_{j}\Lambda_{j}^{-2}U_{j}^{\top}$
@@
$\quad-U_{j\perp}U_{j\perp}^{\top}\zeta^{\top}_{j,\ell,1}U_{j}\Lambda_{j}^{-2}U_{j}^{\top}(\zeta_{j,\ell,1}+\zeta^{\top}_{j,\ell,1})U_{j}\Lambda_{j}^{-2}U_{j}^{\top}$
$\quad+U_{j\perp}U_{j\perp}^{\top}\zeta^{\top}_{j,\ell,1}U_{j}\Lambda_{j}^{-4}U_{j}^{\top}\zeta_{j,\ell,1}U_{j\perp}U_{j\perp}^{\top},$
and $\mathfrak{S}_{j,\ell,3}$ is a remainder term determined by the above equations.
Under $E(C^{\prime}_{2})$, it holds that $\left\lVert Z_{j,\ell}U_{-j}\right\rVert_{2}\leq C^{\prime}_{2}\sigma\sqrt{p}$. Using the fact that $\left\lVert\Lambda_{j}^{-1}G_{j}\right\rVert_{2}=1$ and $\left\lVert\Lambda_{j}^{-1}\right\rVert_{2}\leq\lambda_{\min}^{-1}$, we have
$\left\lVert\mathfrak{S}_{j,\ell,1}\right\rVert_{2}=\left\lVert U_{j}\Lambda_{j}^{-2}G_{j}U_{-j}^{\top}Z_{j,\ell}^{\top}U_{j\perp}U_{j\perp}^{\top}+U_{j\perp}U_{j\perp}^{\top}Z_{j,\ell}U_{-j}G_{j}^{\top}\Lambda_{j}^{-2}U_{j}^{\top}\right\rVert_{2}\lesssim\lambda^{-1}_{\min}\sigma\sqrt{p}.$ (31)
By (27), it holds that
@@
Under $E(C_{2}^{\prime})$, it holds that $\left\lVert\overline{Z}_{j}U_{-j}\right\rVert_{2}\lesssim\sigma\sqrt{p/L}$ and thus
$\left\lVert\overline{\mathfrak{S}_{j,1}}\right\rVert_{2}=\left\lVert U_{j}\Lambda_{j}^{-2}G_{j}U_{-j}^{\top}\overline{Z}^{\top}_{j}U_{j\perp}U_{j\perp}^{\top}+U_{j\perp}U_{j\perp}^{\top}\overline{Z}_{j}U_{-j}G_{j}^{\top}\Lambda_{j}^{-2}U_{j}^{\top}\right\rVert_{2}\lesssim\lambda^{-1}_{\min}\sigma\sqrt{p/L}.$ (34)